OpenStudy (anonymous):

WILL MEDAL AND FAN What is the sum? –25 + (–18)

jimthompson5910 (jim_thompson5910):

You have 2 negatives being added. You can add them up without the negative signs. Then make the final result negative

jimthompson5910 (jim_thompson5910):

example: -2 + (-7) = -9 because 2+7 = 9. Then make the result negative.

jimthompson5910 (jim_thompson5910):

in my example it is, but my example doesn't match your problem

jimthompson5910 (jim_thompson5910):

the answer to your question isn't 9 or -9

OpenStudy (anonymous):

I know.. So is the answer to the problem is -7?

OpenStudy (anonymous):

There ya go lol

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

no you add 25 to 18 to get ??? then you make that result negative

OpenStudy (anonymous):

Oh I did subtraction instead of addition. the real answer is 43?

jimthompson5910 (jim_thompson5910):

-43

jimthompson5910 (jim_thompson5910):

25+18 = 43 make the result negative 43 ----> -43

jimthompson5910 (jim_thompson5910):

you can think of it as (25 ious) + (18 ious) = (43 ious) 43 ious = -43 dollars

OpenStudy (anonymous):

oh damn Ithought it said subtract..